OpenStudy (sleepyjess):

Find the center, vertices, and foci of the ellipse with equation 2x^2 + 6y^2 = 12

OpenStudy (freckles):

have you put it in this form: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1 \]

OpenStudy (sleepyjess):

No, I'm not really sure how to do that :/

OpenStudy (freckles):

also when it is in this form the center is (0,0) :) \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 \text{ has center } (h,k)\]

OpenStudy (freckles):

well do you know that 2=1/(1/2)

OpenStudy (freckles):

\[a=\frac{1}{\frac{1}{a}}\] assuming a isn't 0 of course

OpenStudy (sleepyjess):

Would a be 2?

OpenStudy (freckles):

before we look at that we want 1 on the side that is 12

OpenStudy (freckles):

so divide both sides by 12

OpenStudy (sleepyjess):

1/6x^2 + 1/2y^2 = 1

OpenStudy (freckles):

\[\frac{2}{12} x^2+\frac{6}{12}y^2=1 \\ \\ \text{ now using the one property I mentioned we have } \\ \frac{x^2}{\frac{12}{2}}+\frac{y^2}{\frac{12}{6}}=1 \] and I suppose I could have reduce those fractions before doing that but i will let you do that

OpenStudy (anonymous):

WOW

OpenStudy (sleepyjess):

\(\dfrac{x^2}{6}+\dfrac{y^2}{2}=1\)

OpenStudy (freckles):

peace sleepy :)

OpenStudy (sleepyjess):

@freckles could you help again?

OpenStudy (freckles):

ok so we have a^2=6 and b^2=2

OpenStudy (freckles):

so what is a and b

OpenStudy (sleepyjess):

\(a = \sqrt6\\b=\sqrt2\)

OpenStudy (freckles):

 ugly ellispe but the major axis is the longer one which in this case is the y-axis or also called x=0